A firm can be worth $90 or $310 with equal probability. The firms debt consists of a zero -coupon bond with a face value of $180 that matures at the end of one year. Assume risk neutrality and a cost of capital of 11%. What is the value of this firms equity? O 20 0-23 O 18 O 39
Answer #1

Answer: 39

Expected Value of firm = 90*.5 + 310*.5

= 45 + 155

= 200

Value of Zero Coupon Bond = Face Value / (1+rate)^Time to maturity

= 180/1.11^1

= 162 (rounded to whole number)

Value of firm's Equity = Expected Value of firm - Value of Zero Coupon Bond

= 200-162

= 38